OpenStudy (anonymous):

IS JB GAY Evaluate the expression using the given value of the variable. 2 – 8x + 3x when x = 5 –23 –53 57 –15

OpenStudy (girlmeetsworld):

JB isnt gay. just an fyi

OpenStudy (anonymous):

I LOVE JB

OpenStudy (girlmeetsworld):

me too hes cute.

OpenStudy (anonymous):

NO HES NOT CUTE hes super hot

OpenStudy (anonymous):

ummm idk what that means

OpenStudy (anonymous):

No JB is Hot Af. I fangirl over him in front of my mom yesturday bc he was on tv performing his single 'Sorry'

OpenStudy (anonymous):

The answer is -23 bc when you plug in 5 to the equation you get -23 2-8x+3x 8x=35 3x=15 so now the equation is 2-35+15 2 - 35 =-33 -33+15=-23

OpenStudy (girlmeetsworld):

i like you Ariana pic. :)